it takes kevin 3 hours more than john to paint a fence. together they painted for 3 hours, than kevin left and john finished the job in one additional hour. find how long it would take each boy to paint the fence alone?
Found 2 solutions by rwm, lwsshak3:
Answer by rwm(914)

k=j+3
x/k+x/j=1
3/k+3/j+1/j=1
j = 6, k = 9, x = 18/5=3.6
john would take 6 hrs alone
kevin would take 9 hours alone
together they would take 3.6 hours.

Answer by lwsshak3(11628)

it takes kevin 3 hours more than john to paint a fence. together they painted for 3 hours, than kevin left and john finished the job in one additional hour. find how long it would take each boy to paint the fence alone?
..
let x= hrs John takes to finish the job alone
1/x=John's work rate
x+3 =hrs Kevin takes to finish the job alone
1/(x+3)=Kevin's work rate
..
Kevin worked for 3 hours
John worked for 4 hours
3/(x+3)+4/x)=100%=1
LCD:x(x+3)
3x+4x+12=x(x+3)=x^2+3x
x^2-4x-12=0
(x-6)(x+2)=0
x=6
or
x=-2 (reject,x>0)
..
ans:
John takes 6 hrs to paint the fence alone
Kevin takes 9 hrs to paint the fence alone